OpenStudy (anonymous):

write the polynomial whose zeroes are -1/2 and 2

OpenStudy (lgbasallote):

if zeroes are -1/2 and 2 that means x = -1/2 and x = 2 right? 2x = -1 and x = 2 2x + 1 and x - 2 will then be your factors (2x+1)(x-2) <--expand

OpenStudy (anonymous):

i jst wanna knw d answers r correct

OpenStudy (maheshmeghwal9):

If the roots are a and b then,eq. is \[x^2-(a+b)x+ab=0\] now,solve.

OpenStudy (anonymous):

c mahesh is igbasal's answer correct

OpenStudy (maheshmeghwal9):

i am seeing that. wait!

OpenStudy (anonymous):

k

OpenStudy (maheshmeghwal9):

ya ! it is correct

OpenStudy (anonymous):

k

OpenStudy (maheshmeghwal9):

thus, Answer is\[2x^2-3x-2=0\]

OpenStudy (anonymous):

k